OpenStudy (ksaimouli):

secy sin(pie/2-y) use basic identites and simplyfy? plzzzzzzzzzz help

OpenStudy (ksaimouli):

plz help

OpenStudy (anonymous):

what is secy?????

OpenStudy (anonymous):

\[sec(y)sin\left(\frac{\pi}{2}-y\right)=\frac{1}{cos(y)}\times cos(y)=1\]

OpenStudy (anonymous):

here we know that secy=1/cosy...so we can rewrite the given equation as.. sin(pi/2-y)/cosy... now...we have the basic formula for sin(a-b) as=sina*cosb-cosa*sinb.. by applying that we will get... (sin(90)cosy-cos(90)siny)/cosy.. after plugging the value of sin(90)=1 and cos(90)=0..we will get.. cosy/cosy=1 ans